OpenStudy (anonymous):

use substitution method (1) 0.3x + 0.2y = 5 (2) 0.5x + 0.4y = 11

OpenStudy (anonymous):

(1) 0.3x +0.2y =5 x=(5-0.2y)/0.3 (2) 0.5x + 0.4y = 11 0.5((5-0.2y)/3)+0.4y = 11 (2.5-0.10y)/3 +0.4y = 11 2.5/3 - 0.10/3y + 0.4y = 11 0.4y - 0.10/3y = 11 - 2.5/3 y = (11-2.5/3)/(0.4-0.10/3)

OpenStudy (anonymous):

thank you so much